Determine the mass of one mole (molar mass) of calcium chloride (CaCl2).
1
Expert's answer
2015-12-01T11:06:20-0500
Solution:
The mass of one mole of calcium chloride (CaCl2) is equal to the sum of masses of one mole of Calcium and two moles of Chlorine.
M (CaCl2) = m(Ca) + 2*m(Cl) = 40.078 + 2*35.446 = 110.97 (g)

Answer: 110.97 g.